### v1

@@ Abstract
Moreover, we present an efficient algorithm for calculating an eight-partition of a set of $n$ points in $\mathbb{R}^{3}$ (with prescribed normal direction of one of the planes) in time $O^{*}(n^{5/2})$.
@@ 1 Introduction
Geometric methods for partitioning space, point sets, or other geometric objects are a central topic in discrete and computational geometry. Partitioning results are often proved using topological methods and also play an important role in topological combinatorics [8, 17, 19, 22]. A classical example is the famous Ham-Sandwich Theorem, which goes back to the work of Steinhaus, Banach, Stone, and Tukey (see [19, Sec. 1] for more background and references). A “discrete” version of this theorem asserts that, given any $d$ finite point sets $P_{1},\dots,P_{d}$ in $\mathbb{R}^{d}$, there is an (affine) hyperplane $H$ that simultaneously bisects all $P_{i}$, i.e., each of the two open half-spaces determined by $H$ contains at most $|P_{i}|/2$ points, $1\leq i\leq d$. This follows (by a standard limit argument, see [17, Sec. 3.1]) from the following “continuous” version: Let $\mu_{1},\dots,\mu_{d}$ be mass distributions in $\mathbb{R}^{d}$, i.e., finite measures such that every open set is measurable and every hyperplane has measure zero. Then there exists a hyperplane $H$ such that $\mu_{i}(H^{+})=\mu_{i}(H^{-})=\frac{1}{2}\mu_{i}(\mathbb{R}^{d})$ for $1\leq i\leq d$, where $H^{+}$ and $H^{-}$ are the two open half-spaces bounded by $H$.
@@
Given a mass distribution $\mu$ in $\mathbb{R}^{3}$ and a vector $v\in\mathbb{S}^{2}$, there exists an eight-partition $(H_{1},H_{2},H_{3})$ of $\mu$ such that the intersection of the two planes $H_{1}$ and $H_{2}$ is a line in direction $v$.
@@
We remark that, in general, $d$ hyperplanes in $\mathbb{R}^{d}$ are described by $d^{2}$ independent parameters, while the condition that $2^{d}$ orthants have equal mass can be expressed by $12^{d}-$ 1 equations. For $d=3$, this leaves $9-7=2$ degrees of freedom, which allows for any one of the additional conditions imposed in Theorems 1.1, 1.2, and 1.3, respectively. On the other hand, for $d\geq 5$, we have $d^{2}<2^{d}-1$, so intuitively Grünbaum’s problem is overconstrained. Avis [4] made this precise and constructed explicit counterexamples using the well-known moment curve $\gamma=\{(t,t^{2},\dots,t^{d})\colon t\in\mathbb{R}\}$ in $\mathbb{R}^{d}$. The crucial fact is that any hyperplane intersects the moment curve $\gamma$ in at most $d$ points ([17, Lemma 1.6.4]). Thus, for $d\geq 5$, a mass distribution supported on $\gamma$ admits no $2^{d}$-partition because any $d$ hyperplanes intersect $\gamma$ in at most $d^{2}$ points, which subdivide $\gamma$ into at most $d^{2}+1$ intervals, hence there are always at least $02^{d}-d^{2}-1>$ 0 orthants that do not intersect $\gamma$ and hence contain no mass. The last remaining case $d=4$ of Grünbaum’s problem, i.e., the question whether any mass distribution in $\mathbb{R}^{4}$ admits a $16$-partition by four hyperplanes, remains stubbornly open (see [5], [8, Conjecture 7.2], [17, pp. 50–51], and [19, Problem 2.1.4] for more background and related open problems).
@@
As remarked above, the corresponding problem of computing a four-partition of a planar point can be reduced to finding a Ham-Sandwich cut of two planar point sets that are separated by a line; Megiddo [18] showed that this can be done in linear time.
Computing a Ham-Sandwich cut in $\mathbb{R}^{3}$ can be done efficiently, in time $O^{*}(n^{3/2})$ [16] (where $n$ is the total number of points and the $O^{*}(\cdot)$-notation suppresses polylogarithmic factors). In general, the best known algorithm for computing Ham-Sandwich cuts in fixed dimension $d\geq 3$ runs in time $O(n^{d-1-\alpha_{d}})$ where $\alpha_{d}>0$ a constant depending only on $d$ [16], and a decision variant of the problem becomes computationally hard when the dimension is part of the input, see, e.g., [14]. However, the problem of computing eight-partitions in $\mathbb{R}^{3}$ seems significantly more difficult, and there is no known way of reducing it to the computation of a Ham-Sandwich cut; in particular, given two planes $H_{1}$ and $H_{2}$ that four-sect a finite point set $P$ (in the sense that every one of the four open orthants determined by $H_{1}$ and $H_{2}$ contains at most $|P|/4$ points), there generally need not exist a third plane $H_{3}$ such that $H_{1},H_{2},H_{3}$ form an eight-partition.
The following concept is useful for characterizing the complexity of Problem 1.4. A halving plane for an $n$-point set for $n$ odd in $\mathbb{R}^{3}$ in general position is a plane that passes through three of the points and contains exactly $(n-3)/2$ points on each side. Let $h_{3}(n)$ be the maximum number of halving planes for an $n$-point set $\mathbb{R}^{3}$ as above. The best known upper and lower bounds for $h_{3}(n)$ are $O(n^{5/2})$ and $\Omega(n^{2}e^{\sqrt{\log n}})$, due to Sharir, Smorodinsky, and Tardos [20] and Tóth [21], respectively. A brute-force algorithm that checks every triple of halving planes solves Problem 1.4 in time roughly proportional to $(h_{3}(n))^{3}$.
Yao et al. [23] and Edelsbrunner [9] gave a $O(n^{6})$ time algorithm for Problem 1.4 that computes an eight-partition (with a prescribed normal direction for one of the planes, as in Theorem 1.1) by an expensive search, using the fact that only two planes need to be identified. Fixing one plane and performing a brute-force search for the remaining two would yield an algorithm with a running time comparable to $(h_{3}(n))^{2}$.
Here, we present, to our knowledge, the fastest known algorithm for Problem 1.4. Roughly speaking, our algorithm runs in time near-linear in $h_{3}(n)$ rather than quadratic in it.
@@
An eight-partition of $n$ points in general position in $\mathbb{R}^{3}$, with a prescribed normal vector for one of the planes, can be computed in time $O^{*}(n^{5/2})$; the $O^{*}(\cdot)$-notation suppresses polylogarithmic factors.
Our algorithm can be seen as a constructive version of Hadwiger’s proof [12]. We start by bisecting the point set by a plane with a fixed normal direction, which partitions the initial point set into two subsets of “red” and “blue” points, respectively, of equal size. After that, our algorithm finds two more planes that simultaneously four-sect both the red points and the blue points.
It remains an open question whether Theorem 1.2 or our own Theorem 1.3 can also be used to obtain an efficient algorithm for Problem 1.4. It would also be interesting to decide whether there is an algorithm for Problem 1.4 with running time $o(h_{3}(n))$.
@@ 2.1 Notation and preliminaries
In what follows, it will often be convenient to assume that the mass distributions we work with have connected support where the support of a mass distribution $\mu$ is $\operatorname{Supp}(\mu)\coloneqq\{x\in\mathbb{R}^{3}:\mu(B_{r}(x))>0\text{ for every }r>0\}$ and $B_{r}(x)$ denotes the ball of radius $r>0$ centred at $x$. By a standard limit argument (see Lemma A.5), the existence of eight-partitions for mass distributions with connected support implies the existence of eight-partitions for the general case. Hereafter, unless stated otherwise, we assume, without loss of generality, that every mass distribution has connected support.
@@
For $\alpha=(\alpha_{1},\dots,\alpha_{k})\in\mathbb{Z}_{2}^{k}=\{+,-\}^{k}$, we define the open orthant determined by $\mathcal{H}$ and $\alpha$ as $\mathcal{O}_{\alpha}^{\mathcal{H}}:=H_{1}^{\alpha_{1}}\cap\dots\cap H_{k}^{\alpha_{k}}$. Given a mass distribution $\mu$ in $\mathbb{R}^{3}$, we say that an ordered $k$-tuple $\mathcal{H}=(H_{1},\dots,H_{k})$ of planes ($k\leq 3$) forms a $2^{k}$-partition of $\mu$ if every orthant contains $1/2^{k}$ of the mass, i.e., $\mu(\mathcal{O}_{\alpha}^{\mathcal{H}})=\mu(\mathbb{R}^{3})/2^{k}$ for every $\alpha\in\{+,-\}^{k}$. For $k=1,2,3$, this corresponds to the notions of bisecting, four-secting, and eight-partitioning $\mu$ as mentioned in the introduction. Analogously, we say that $\mathcal{H}$ forms a $2^{k}$-partition of a finite point set $P$ in $\mathbb{R}^{3}$ if $\lvert P\cap\mathcal{O}^{\mathcal{H}}_{\alpha}\rvert\leq\frac{\lvert P\rvert}{2^{k}}$ for all $\alpha$.
@@
where $p(\alpha,\beta)$ is the number of coordinates where both $\alpha$ and $\beta$ are --. As an example, with $\mathcal{H}:=\mathcal{H}_{v}=(H_{1},H_{2},H_{3})$ and $\alpha=--+\in\mathbb{Z}_{2}^{3}\setminus\{\bm{+}\}$, we obtain
@@
    Let $\alpha\in\mathbb{Z}_{2}^{k-1}\setminus\{\bm{+}\}$ and let $\mathcal{H}=(H_{1},\dots,H_{k})$ be a $k$-tuple of planes, then $F_{+\alpha}(\mathcal{H})=F_{\alpha}((H_{2},\dots,H_{k}))$ (the equivalent statement holds for any other entry of a $k$-tuple $\alpha_{1}\cdots\alpha_{k}$ instead of just for $\alpha_{1}$).
@@
$\mu(\mathcal{O}^{\mathcal{H}}_{\beta})=\mu(\mathcal{O}^{\mathcal{H}}_{+\beta})+\mu(\mathcal{O}^{\mathcal{H}}_{-\beta}).$
From the definition of $F_{\tilde{\alpha}}$, if $\tilde{\alpha}=+\alpha$ then for any $\beta\in\mathbb{Z}_{2}^{k-1}$ the two orthants $\mathcal{O}^{\mathcal{H}}_{+\beta}$ and $\mathcal{O}^{\mathcal{H}}_{-\beta}$ are counted with the same sign in the sum, therefore
@@
Suppose first that $k=1$ and that $\mathcal{H}=H$. The only alternating sum is $F_{-}(H)=\mu(H^{+})-\mu(H^{-})$ and $F_{-}(H)=0$ implies that $H$ bisects.
@@ 2.2 The main topological result
Our goal is to prove the following result — a more precise statement of Theorem 1.3:
@@
Given a mass distribution $\mu$ and a direction $p\in\mathbb{S}^{2}$, there exists a triple $\mathcal{H}=(H_{1},H_{2},H_{3})$ of oriented planes that eight-partition $\mu$ so that the oriented direction of the intersection $H_{1}\cap H_{2}$ is $p$.
By Lemma A.5, it is sufficient to prove Theorem 2.2 for mass distributions with connected support. We require the following technical lemma about partitioning a mass distribution on $\mathbb{R}^{2}$, due to Blagojević and Karasev [6]. For completeness, the proof is given in Appendix C. {restatable}[Four-partitioning a mass distribution in $\mathbb{R}^{2}$ [6]]lemmablagkarasev Let $\mu^{\#}$ be a mass distribution (with connected support) on $\mathbb{R}^{2}$ and $v\in\mathbb{S}^{1}$. Then there exists a pair $(\ell_{1},\ell_{2})$ of lines in $\mathbb{R}^{2}$ that four-partitions $\mu^{\#}$ and such that $v$ bisects the angle between $\ell_{1}$ and $\ell_{2}$.
@@
Without loss of generality, let $v=(0,0,1)$. Our proof proceeds in two steps. In the first step, we construct a map $\Phi\colon\mathbb{S}^{1}\times\mathbb{S}^{3}\rightarrow\mathbb{R}^{4}$ whose zeros codify equipartitions of $\mu$; then we prove that $\Phi$ is equivariant with respect to a suitable choice of actions of $G:=\mathbb{Z}_{4}\times\mathbb{Z}_{2}$ on the two spaces. In the second step we show that any continuous $G$-equivariant map $\Psi\colon\mathbb{S}^{1}\times\mathbb{S}^{3}\rightarrow\mathbb{R}^{4}$ has to have a zero.
Step 1: The key step in constructing the map $\Phi$ is to show that we can parameterize pairs of planes that have intersection direction $p$ and four-sect $\mu$, by a vector in $\mathbb{S}^{1}$.
We project $\mu$ to the $xy$-plane to obtain a mass distribution $\mu^{\#}$ on $\mathbb{R}^{2}$. Lemma 2.2 guarantees that, once we fix a direction $v\in\mathbb{S}^{1}\hookrightarrow\mathbb{S}^{2}$ (inclusion as the horizontal equator in $\mathbb{S}^{2}$) there are two lines in the $xy$-plane $\ell_{1}=\mathbb{R}\vec{\ell}_{1}(v)+a_{0}(v)$ and $\ell_{2}=\mathbb{R}\vec{\ell}_{2}(v)+a_{2}(v)$ that four-sect the projected measure $\mu^{\#}$. Define $H_{i}(v)\coloneqq(h_{i}(v),a_{i}(v))$ to be the (oriented) span of $\ell_{i}(v)$ and $v$; the two planes now four-sect $\mu$ and have the desired intersection direction.
Now let $g_{1}$ be a generator of $\mathbb{Z}_{4}\times\{\bm{+}\}\subseteq G$ and define its action on $\mathbb{S}^{1}$ by a counterclockwise rotation by $\frac{\pi}{2}$. We use $g_{1}\cdot v$ to denote the action of $g_{1}$ on $v$. Then, by the uniqueness in Lemma 2.2, we have that
@@
Intuitively, if we consider the planar problem with the bisecting vector $v$ rotated by $\frac{\pi}{2}$, by uniqueness the affine lines that split the measure are the same. However, while the direction chosen as $\vec{\ell}_{1}$ in the rotated problem is the direction $\vec{\ell}_{2}$ in the previous configuration, the “rotated” $\vec{\ell}_{2}$ is $-\vec{\ell}_{1}$ in the original problem (see Figure 1).
@@
Using this construction, we can define a function $\mathbb{S}^{1}\rightarrow\mathbb{S}^{3}\times\mathbb{S}^{3}$ by $v\mapsto(H_{1}(v),H_{2}(v))$. It follows from eq. (1) that $g_{1}\cdot v$ is mapped to $(H_{2}(v),-H_{1}(v))$, therefore, if we fix the corresponding action [^1] Formally, for any $(x,y)\in\mathbb{S}^{3}\times\mathbb{S}^{3}$ the generator $g_{1}$ of $\mathbb{Z}_{4}\times\{+\}\subseteq G$ acts by $g_{1}\cdot(x,y)=(y,-x)$. of $\mathbb{Z}_{4}$ on $\mathbb{S}^{3}\times\mathbb{S}^{3}$, the map is $\mathbb{Z}_{4}$-equivariant.
The group $\{e\}\times\mathbb{Z}_{2}$ acts by antipodality on $\mathbb{S}^{3}$; therefore, if $G$ acts on $\left(\mathbb{S}^{3}\times\mathbb{S}^{3}\right)\times\mathbb{S}^{3}$ component-wise, the map $\Phi\colon\mathbb{S}^{1}\times\mathbb{S}^{3}\to\left(\mathbb{S}^{3}\times\mathbb{S}^{3}\right)\times\mathbb{S}^{3}$ defined as $\Phi(v,w)\coloneqq(H_{1}(v),H_{2}(v),w)$ is $G$-equivariant.
@@
To compute the action of $G$ on the alternating sums, it is enough to specify what happens on $g_{1}$ (a generator of $\mathbb{Z}_{4}\times\{e\}$) and $g_{2}$ (a generator of $\{e\}\times\mathbb{Z}_{2}$). If we act with $g_{1}$, we obtain
@@
is $G$-equivariant. By Observation 2.1, the zeros of $\Psi$ are exactly the configurations of planes that eight-partition the measure and have the desired intersection property.
@@
Denote by $\Psi_{a}$, for $a\in\mathbb{S}^{1}$, the map $\Psi_{a}\colon\mathbb{S}^{3}\rightarrow\mathbb{S}^{3}$, $\Psi_{a}(p)=\overline{\Psi}(a,p)$; this function has two key properties:

### v4

@@ Abstract
Moreover, we present an efficient algorithm for calculating an eight-partition of a set of $n$ points in $\mathbb{R}^{3}$ (with prescribed normal direction of one of the planes) in time $O^{*}(n^{7/3})$.
@@ 1 Introduction
Geometric methods for partitioning space, point sets, or other geometric objects are a central topic in discrete and computational geometry. Partitioning results are often proved using topological methods and also play an important role in topological combinatorics [8, 19, 21, 24]. A classical example is the famous Ham-Sandwich Theorem, which goes back to the work of Steinhaus, Banach, Stone, and Tukey (see [21, Sec. 1] for more background and references). A “discrete” version of this theorem asserts that, given any $d$ finite point sets $P_{1},\dots,P_{d}$ in $\mathbb{R}^{d}$, there is an (affine) hyperplane $H$ that simultaneously bisects all $P_{i}$, i.e., each of the two open half-spaces determined by $H$ contains at most $|P_{i}|/2$ points, $1\leq i\leq d$. This follows (by a standard limit argument, see [19, Sec. 3.1]) from the following “continuous” version: Let $\mu_{1},\dots,\mu_{d}$ be mass distributions in $\mathbb{R}^{d}$, i.e., finite measures such that every open set is measurable and every hyperplane has measure zero. Then there exists a hyperplane $H$ such that $\mu_{i}(H^{+})=\mu_{i}(H^{-})=\frac{1}{2}\mu_{i}(\mathbb{R}^{d})$ for $1\leq i\leq d$, where $H^{+}$ and $H^{-}$ are the two open half-spaces bounded by $H$.
@@
Given a mass distribution $\mu$ in $\mathbb{R}^{3}$ and a vector $v\in\mathbb{S}^{2}$, there exists an eight-partition $(H_{1},H_{2},H_{3})$ of $\mu$ such that the intersection of the two planes $H_{1}$ and $H_{2}$ is a line in the direction of $v$.
@@
We remark that, in general, $d$ hyperplanes in $\mathbb{R}^{d}$ are described by $d^{2}$ independent parameters, while the condition that $2^{d}$ orthants have equal mass can be expressed by $12^{d}-$ 1 equations. For $d=3$, this leaves $9-7=2$ degrees of freedom, which allows for any one of the additional conditions imposed in Theorems 1.1, 1.2, and 1.3, respectively. On the other hand, for $d\geq 5$, we have $d^{2}<2^{d}-1$, so intuitively Grünbaum’s problem is overconstrained. Avis [4] made this precise and constructed explicit counterexamples using the well-known moment curve $\gamma=\{(t,t^{2},\dots,t^{d})\colon t\in\mathbb{R}\}$ in $\mathbb{R}^{d}$. The crucial fact is that any hyperplane intersects the moment curve $\gamma$ in at most $d$ points ([19, Lemma 1.6.4]). Thus, for $d\geq 5$, a mass distribution supported on $\gamma$ admits no $2^{d}$-partition because any $d$ hyperplanes intersect $\gamma$ in at most $d^{2}$ points, which subdivide $\gamma$ into at most $d^{2}+1$ intervals, hence there are always at least $02^{d}-d^{2}-1>$ 0 orthants that do not intersect $\gamma$ and hence contain no mass. The last remaining case $d=4$ of Grünbaum’s problem, i.e., the question whether any mass distribution in $\mathbb{R}^{4}$ admits a $16$-partition by four hyperplanes, remains stubbornly open (see [5], [8, Conjecture 7.2], [19, pp. 50–51], and [21, Problem 2.1.4] for more background and related open problems).
@@
As noted above, the corresponding problem of computing a four-partition of a planar point set can be reduced to finding a Ham-Sandwich cut of two planar point sets that are separated by a line. Megiddo [20] showed that this can be done in linear time.
To characterize the complexity of Problem 1.4, we introduce the following concept. A halving line (resp., halving plane) for an $n$-point set in $\mathbb{R}^{2}$ (resp., $\mathbb{R}^{3}$) in general position is a line (resp., plane) that passes through two (resp., three) of the points and divides the remaining ones as equally as possible. Let $h_{2}(n)$ (resp., $h_{3}(n)$) denote the maximum number of halving lines (resp., planes) for an $n$-point set in $\mathbb{R}^{2}$ (resp., $\mathbb{R}^{3}$). The best known upper and lower bounds for $h_{2}(n)$ are $O(n^{4/3})$, due to Dey [9], and $\Omega(ne^{\sqrt{\log n}})$, due to Tóth [23], respectively. For $h_{3}(n)$, the best-known bounds are are $O(n^{5/2})$, due to Sharir, Smorodinsky, and Tardos [22], and $\Omega(n^{2}e^{\sqrt{\log n}})$, due to Tóth [23].
By a result of Lo, Matoušek, and Steiger [18, Proposition 2], a Ham-Sandwich cut of $n$ points in $\mathbb{R}^{3}$ can be computed in time $O^{*}(h_{2}(n))=O^{*}(n^{4/3})$ (see also [14]), where the $O^{*}(\cdot)$-notation suppresses polylogarithmic factors. However, computing eight-partitions in $\mathbb{R}^{3}$ appears to be significantly more difficult. Unlike the planar four-partition problem, there is no known way of reducing it to the computation of a Ham-Sandwich cut. In particular, given two planes $H_{1}$ and $H_{2}$ that four-partition a finite point set $P$ in $\mathbb{R}^{3}$ (in the sense that every one of the four open quadrants determined by $H_{1}$ and $H_{2}$ contains at most $|P|/4$ points), there generally need not exist a third plane $H_{3}$ such that $H_{1},H_{2},H_{3}$ form an eight-partition.
We note that, for fixed dimension $d\geq 3$, the best known algorithm for computing Ham-Sandwich cuts in $\mathbb{R}^{d}$ runs in time $O(n^{d-1-\alpha_{d}})$ where $\alpha_{d}>0$ is a constant depending only on $d$ [18]. When the dimension is part of the input, a decision variant of the problem becomes computationally hard, see, e.g., [16].
A brute-force algorithm that checks every triple of halving planes solves Problem 1.4 in time comparable to $O(h_{3}(n)^{3})=O(n^{15/2})$. Yao et al. [25] and Edelsbrunner [10] gave a $O(n^{6})$-time algorithm that computes an eight-partition (with a prescribed normal direction for one of the planes, as in Theorem 1.1) by an exhaustive search, using the fact that only two planes need to be identified. Fixing one plane and performing a brute-force search for the remaining two would yield an algorithm with a running time comparable to $O(h_{3}(n)^{2})=O(n^{5})$.
Here, we present, to our knowledge, the fastest known algorithm for Problem 1.4. Roughly speaking, our algorithm runs in time near-linear in $h_{3}(n)$ rather than quadratic in it. Slightly more precisely, our algorithm runs in time near-linear in $nh_{2}(n)$, which is not known to be $o(h_{3}(n))$, but for which the best known upper bound is strictly stronger; see Theorem 4.2 and Fact 2:
@@
An eight-partition of $n$ points in general position in $\mathbb{R}^{3}$, with a prescribed normal vector for one of the planes, can be computed in time $O^{*}(nh_{2}(n))$, hence $O^{*}(n^{7/3})$; here, the $O^{*}(\cdot)$-notation suppresses polylogarithmic factors.
Our algorithm can be seen as a constructive version of Hadwiger’s proof [13]. We start by bisecting the point set by a plane with a fixed normal direction, which partitions the initial point set into two subsets of “red” and “blue” points, respectively, of equal size. After that, our algorithm finds two more planes that simultaneously four-partition both the red and the blue points.
It remains an open question whether Theorem 1.2 or our own Theorem 1.3 can also be used to obtain an efficient algorithm for Problem 1.4. It would also be interesting to decide whether there is an algorithm for Problem 1.4 with running time $o(nh_{2}(n))$.
@@ 2.1 Notation and preliminaries
In what follows, it will often be convenient to assume that the mass distributions we work with have connected support, where the support of a mass distribution $\mu$ is $\operatorname{Supp}(\mu)\coloneqq\{x\in\mathbb{R}^{3}:\mu(B_{r}(x))>0\text{ for every }r>0\}$ and $B_{r}(x)$ denotes the ball of radius $r>0$ centered at $x$.
By a standard limit argument (see Lemma A.5), the existence of eight-partitions for mass distributions with connected support implies the existence of eight-partitions for the general case. Hereafter, unless stated otherwise, we assume, without loss of generality, that every mass distribution has connected support.
@@
For $\alpha=(\alpha_{1},\dots,\alpha_{k})\in\mathbb{Z}_{2}^{k}=\{+,-\}^{k}$, we define the open orthant determined by $\mathcal{H}$ and $\alpha$ as $\mathcal{O}_{\alpha}^{\mathcal{H}}:=H_{1}^{\alpha_{1}}\cap\dots\cap H_{k}^{\alpha_{k}}$. Given a mass distribution $\mu$ in $\mathbb{R}^{3}$, we say that an ordered $k$-tuple $\mathcal{H}=(H_{1},\dots,H_{k})$ of planes ($k\leq 3$) forms a $2^{k}$-partition of $\mu$ if every orthant contains $1/2^{k}$ of the mass, i.e., $\mu(\mathcal{O}_{\alpha}^{\mathcal{H}})=\mu(\mathbb{R}^{3})/2^{k}$ for every $\alpha\in\{+,-\}^{k}$. For $k=1,2,3$, this corresponds to the notions of bisecting, four-partitioning, and eight-partitioning $\mu$ as mentioned in the introduction. Analogously, we say that $\mathcal{H}$ forms a $2^{k}$-partition of a finite point set $P$ in $\mathbb{R}^{3}$ if $\lvert P\cap\mathcal{O}^{\mathcal{H}}_{\alpha}\rvert\leq\frac{\lvert P\rvert}{2^{k}}$ for all $\alpha$.
@@
where $p(\alpha,\beta)$ is the number of coordinates where both $\alpha$ and $\beta$ are --. The functions $F_{\alpha}$ were also utilized in the proof of Theorem 1.1 in [25].
As an example, with $\mathcal{H}:=\mathcal{H}_{v}=(H_{1},H_{2},H_{3})$ and $\alpha=--+\in\mathbb{Z}_{2}^{3}\setminus\{\bm{+}\}$, we obtain
@@
    Let $\alpha\in\mathbb{Z}_{2}^{k-1}\setminus\{\bm{+}\}$ and let $\mathcal{H}=(H_{1},\dots,H_{k})$ be a $k$-tuple of planes. Then $F_{+\alpha}(\mathcal{H})=F_{\alpha}((H_{2},\dots,H_{k}))$ (the equivalent statement holds for any other entry of a $k$-tuple $(\alpha_{1},\cdots,\alpha_{k})$ instead of just for $\alpha_{1}$).
@@
$\mu(\mathcal{O}^{(H_{2},\dots,H_{k})}_{\beta})==\mu(\mathcal{O}^{\mathcal{H}}_{+\beta})+\mu(\mathcal{O}^{\mathcal{H}}_{-\beta}).$
If $\tilde{\alpha}=+\alpha$ then, from the definition of $F_{\tilde{\alpha}}$, for any $\beta\in\mathbb{Z}_{2}^{k-1}$ the two orthants $\mathcal{O}^{\mathcal{H}}_{+\beta}$ and $\mathcal{O}^{\mathcal{H}}_{-\beta}$ are counted with the same sign in the sum, therefore
@@
Suppose first that $k=1$ and that $\mathcal{H}=H$. The only alternating sum is $F_{-}(H)=\mu(H^{+})-\mu(H^{-})$ and $F_{-}(H)=0$ implies that $H$ bisects $\mu$.
@@ 2.2 The main topological result
Our goal is to prove the following result, which is a more precise statement of Theorem 1.3:
@@
Given a mass distribution $\mu$ and a direction $u\in\mathbb{S}^{2}$, there exists a triple $\mathcal{H}=(H_{1},H_{2},H_{3})$ of oriented planes that eight-partition $\mu$ so that the oriented direction of the intersection $H_{1}\cap H_{2}$ is $u$.
By Lemma A.5, it is sufficient to prove Theorem 2.2 for mass distributions with connected support. We require the following technical lemma about partitioning a mass distribution on $\mathbb{R}^{2}$, due to Blagojević and Karasev [6]. For completeness, the proof is given in Appendix B. {restatable}[Four-partitioning a mass distribution in $\mathbb{R}^{2}$ [6]]lemmablagkarasev Let $\mu^{\#}$ be a mass distribution (with connected support) on $\mathbb{R}^{2}$ and $v\in\mathbb{S}^{1}$. Then there exists a pair $(\ell_{1},\ell_{2})$ of lines in $\mathbb{R}^{2}$ that four-partitions $\mu^{\#}$ and such that $v$ bisects the angle between $\ell_{1}$ and $\ell_{2}$.
@@
Without loss of generality, let $u=(0,0,1)$. Our proof proceeds in two steps. In the first step, we construct a map $\Phi\colon\mathbb{S}^{1}\times\mathbb{S}^{3}\rightarrow\mathbb{R}^{4}$ whose zeros codify eight-partitions of $\mu$; then we prove that $\Phi$ is equivariant with respect to a suitable choice of actions of $G:=\mathbb{Z}_{4}\times\mathbb{Z}_{2}$ on the two spaces. In the second step we show that any continuous $G$-equivariant map $\Psi\colon\mathbb{S}^{1}\times\mathbb{S}^{3}\rightarrow\mathbb{R}^{4}$ has to have a zero.
Step 1: The key step in constructing the map $\Phi$ is to show that we can parameterize pairs of planes that have intersection direction $u$ and four-partition $\mu$, by a vector in $\mathbb{S}^{1}$.
We project $\mu$ onto $u^{\bot}$, the plane orthogonal to $u$, to obtain a mass distribution $\mu^{\#}$ on $\mathbb{R}^{2}$. Specifically, identifying the plane with $\mathbb{R}^{2}$, let $A\subseteq\mathbb{R}^{2}$ and set $A\times\mathbb{R}$ to be the cylinder over $A$ in the $u$-direction. Then $\mu^{\#}(A)=\mu(A\times\mathbb{R})$.
Let $v\in\mathbb{S}^{1}\subseteq\mathbb{R}^{2}$. By Lemma 2.2, there are two oriented lines $\ell_{1}(v)$ and $\ell_{2}(v)$ (that we can interpret as points $(\ell^{i}_{1},\ell^{i}_{2},a^{i})\in\mathbb{S}^{2}$) in the plane $u^{\bot}$ such that $v$ bisects the angle between the two and $\ell_{1},\ell_{2}$ four-partition the projected measure $\mu^{\#}$. Define $H_{i}(v)\coloneqq(\ell_{1}^{i}(v),\ell^{i}_{2}(v),0,a^{i}(v))\in\mathbb{S}^{3}$ to be $\ell_{i}(v)\times\mathbb{R}$ the (oriented) span of $\ell_{i}(v)$ and $u$; the two planes now four-partition $\mu$ and have the desired intersection direction.
Now let $g_{1}$ be a generator of $\mathbb{Z}_{4}\times\{\bm{+}\}\subseteq G$ and define its action on $\mathbb{S}^{1}$ by a counterclockwise rotation by $\frac{\pi}{2}$. We use $g_{1}\cdot v$ to denote the action of $g_{1}$ on $v$. Then, by the uniqueness in Lemma 2.2, we have that (see Figure 1):
@@
Using this construction, we can define a function $\mathbb{S}^{1}\rightarrow\mathbb{S}^{3}\times\mathbb{S}^{3}$ by $v\mapsto(H_{1}(v),H_{2}(v))$. It follows from eq. (1) that $g_{1}\cdot v$ is mapped to $(H_{2}(v),-H_{1}(v))$. Therefore, if we fix the corresponding action [^1] Formally, for any $(x,y)\in\mathbb{S}^{3}\times\mathbb{S}^{3}$ the generator $g_{1}$ of $\mathbb{Z}_{4}\times\{+\}\subseteq G$ acts by $g_{1}\cdot(x,y)=(y,-x)$. of $\mathbb{Z}_{4}$ on $\mathbb{S}^{3}\times\mathbb{S}^{3}$, the map is $\mathbb{Z}_{4}$-equivariant.
The group $\{e\}\times\mathbb{Z}_{2}$ acts by antipodality on $\mathbb{S}^{3}$; therefore, if $G$ acts on $\left(\mathbb{S}^{3}\times\mathbb{S}^{3}\right)\times\mathbb{S}^{3}$ component-wise, the map $\Phi\colon\mathbb{S}^{1}\times\mathbb{S}^{3}\to\left(\mathbb{S}^{3}\times\mathbb{S}^{3}\right)\times\mathbb{S}^{3}$ defined as $\Phi(v,w)\coloneqq(H_{1}(v),H_{2}(v),H_{w})$ is $G$-equivariant.
@@
To compute the action of $G$ on the alternating sums, it is enough to specify what happens on $g_{1}$ (a generator of $\mathbb{Z}_{4}\times\{e\}$) and $g_{2}$ (the generator of $\{e\}\times\mathbb{Z}_{2}$). Recalling that $g_{1}\cdot\Phi(v,w)=(H_{2}(v),-H_{1}(v),H_{w})$ and applying Observation 2.1(1), we obtain
$~{}~{}~{}~{}F_{+--}(g_{1}\cdot\Phi(v,w))=F_{+--}((H_{2}(v),-H_{1}(v),H_{w}))=F_{--}((-H_{1}(v),H_{w}))$
$=\mu(-H_{1}(v)^{+}\cap H_{w}^{+})+\mu(-H_{1}(v)^{-}\cap H_{w}^{-})-\mu(-H_{1}(v)^{-}\cap H_{w}^{+})-\mu(-H_{1}(v)^{+}\cap H_{w}^{-})$
$=\mu(H_{1}(v)^{-}\cap H_{w}^{+})+\mu(H_{1}(v)^{+}\cap H_{w}^{-})-\mu(H_{1}(v)^{+}\cap H_{w}^{+})-\mu(H_{1}(v)^{-}\cap H_{w}^{-})$
$=-F_{-+-}(\Phi(v,w))$
Similar computations imply that, if we act with $g_{1}$, we obtain
@@
is $G$-equivariant, i.e., $\Psi(g_{1}\cdot v,w)=g_{1}\cdot\Psi(v,w)$ and $\Psi(v,-w)=-\Psi(v,w)$. By Observation 2.1, the zeros of $\Psi$ are exactly the configurations of planes that eight-partition the measure and have the desired intersection property.
@@
Denote by $\Psi_{a}$, for $a\in\mathbb{S}^{1}$, the map $\Psi_{a}\colon\mathbb{S}^{3}\rightarrow\mathbb{S}^{3}$, $\Psi_{a}(w)=\overline{\Psi}(a,w)$; this function has two key properties: